On average, there are 17 dandelions for every 5 square metres in a large field.
a) If you select a random 5 square metre section of this field, what is the probability that you will find 10
dandelions?
b) If you select a random 3 square metre section of this field, what is the probability that you will find 8
dandelions?
Answer #1

a) For Poisson distribution,

P(X=x)=\frac{e^{-\lambda}\lambda^{x}}{x!}

Given,

\lambda=17

x=10

P(X=10)=\frac{e^{-17}*17^{10}}{10!}

P(X=10)=0.0230

b) For 1 square metre, average number of dandelions= 17/5 =3.4

For 3 square metres, average number of dandelions=(3*3.4)=10.2

  \lambda=10.2

  x=8

P(X=8)=\frac{e^{-10.2}*(10.2)^{8}}{8!}

P(X=8)=0.1080